Grade 11General Physics

On an average, against 10cm of Hg pressure humam heart (per beat) pumps 50cm^3 blood. If in one minute heart beats 72 times, what is the power of heart?

Amit Kumar
8 Years agoGrade 11

1 Answer

Khimraj
8 Years ago
Pressure = 10cm of Hg = 10*1333.2N/m2 = 13332N/m2
\DeltaV = 50cm3 = 50*10-6m3
work done per beat = P\DeltaV =13332*50*10-6J = 6.7J
Total work done = 72*6.7 =482.4J
Power = W/T = 482.4/60 = 8.04W.
Hope it clears. If you like anserthen please approve it